OpenStudy (amurciano2):

How would I integrate the integral of xln(x+1)dx so far I have the u and the dv and I have I=ln(1+x)X^2/2- The integral of x^2/2+2x

OpenStudy (phi):

let u= (x+1) for the integral

OpenStudy (amurciano2):

okay. thank you

OpenStudy (phi):

specifically \[ -\frac{1}{2} \int \frac{x^2}{x+1} \ dx \]